# AI4Research: A Survey of Artificial Intelligence for Scientific Research

### v1

@@ 9.3 AI for Scientific Discovery
Table 8: Representative AI tools and their role in facilitating scientific discovery, with a particular focus on experimental conduction.
@@
requires the collection of scientific evidence, theoretical verification, and theorem proving. Specifically, FV-Generalization Benchmark [586], SCitance [17], and MissciPlus [240] are designed to complete scientific evidence collection. TheoremExplainBench [390] , XClaimCheck [363], SciNews [91], ClaimReview2024+ [75], FactKG [376], TrendFact [900] provide datasets and benchmarks for scientific verification analysis. MiniF2F [925], FIMO [480], MUSTARDSAUCE [323] are used to fine-tune and evaluate LLMs on scientific theorem proving tasks. Furthermore, datasets and benchmarks have expanded into broader domains, including biomedicine [809].
@@
To evaluate model performance in realistic research environments, MLAgentBench [320], Exp-Bench [383], MLRC-Bench [911], MLE-Bench [95], DS-Bench [350], ScienceBoard [727], AutoReproduce [921], SciReplicate-Bench [839], DO Challenge [697], and MLR-Bench [102] assess AI agents’ abilities to perform typical research tasks, such as optimizing CIFAR-10 classifiers and tuning BabyLM. In a similar vein, Hu et al. [307] develop InfiAgent-DABench, which is based on real-world CSV datasets and evaluates models’ ability to interact with tools in end-to-end data analysis tasks. MLGym-Bench [559] is the first Gym environment for machine learning tasks, enabling research on reinforcement learning algorithms for training such agents. ResearchCodeBench [310] enables continuous understanding and advancement of LLM-driven innovation in research code generation. AutoBio [410] is designed to evaluate robotic automation in biology laboratory environments.
@@ 9.4 AI for Academic Writing
The field of AI in academic writing is supported by a comprehensive array of meticulously curated datasets that address various aspects of the academic writing process.
@@ 9.4.1 Semi-Automatic Academic Writing
Several datasets, including FigGen [643], Figuring out Figures [90], SciCapenter [300], and TikZero [55], support figure and formula generation, from text-to-figure creation to automated TikZ code generation. For citation management, datasets like CITEWORTH [822], CiteBART [92], and ScholarCopilot [808] enhance context-aware automatic citation generation [106]. Additionally, FutureGen [35] extracts future work statements from thousands of papers, using LLMs to identify and validate forward-looking scientific content.
@@ 9.5 AI for Academic Peer Reviewing
Research on AI for Academic Peer Reviewing is grounded in diverse datasets, addressing tasks from AI text detection to review generation, quality assessment, and decision support [934, 194, 632, 137, 136, 712, 48]. To simulate realistic peer review interactions, datasets such as PeerRead [358], SPOT [700], NLPeer [180], ReviewMT [739], MOPRD [470], OpenReviewer [327], MASSW [901], COMPARE [690] , PeerArg [720], Re2 [890], ReviewEval [380], AAAR-1 [505], Papereval [321], ORB [733] and ORSUM [886] collect extensive paper and review data from leading conferences or journals, enabling the training and evaluation of LLMs in multi-turn, long-context, or role-based peer reviews [598]. Additionally, LLMart [461] offers a toolkit for evaluating LLM robustness through adversarial testing and prompt optimization, ensuring AI reliability in sensitive academic contexts. To assess LLM review quality more precisely, Shin et al. [679] and Couto et al. [147] analyze the quality of peer review content across multiple predefined aspects, highlighting discrepancies between LLM and human review focus.
@@ 10.4 Explainability and Transparency of AI4Research
As AI models increasingly drive scientific discovery, ensuring their transparency and explainability is essential. Future work should strengthen model interpretability so that researchers can trace how conclusions and recommendations are generated, particularly in high-stakes scientific applications [211, 185]. Efforts to improve explainability fall into two main categories: (1) White-box Analysis: This approach investigates the model’s internal structure by linking specific network “circuits” to conceptual representations. It has attracted considerable interest from both the security and transparency communities [57, 916, 627]. (2) Black-box Analysis: More recent work focuses on interpreting models without direct access to internal parameters. By examining reasoning trajectories and aggregate behavior, black-box methods provide insights into a model’s knowledge representation and enable more reliable control over its outputs [85, 280, 107, 109, 111].

### v2

@@ 9.3 AI for Scientific Discovery
To facilitate scientific discovery, a variety of AI datasets and tools have been developed to assist researchers in different stages of the research process. These tools range from experiment design and management to full-automatic discovery, as shown in Table 8.
@@
Table 8: Representative AI tools and their role in facilitating scientific discovery, with a particular focus on experiment conduction.
@@
requires the collection of scientific evidence, theoretical verification, and theorem proving. Specifically, FV-Generalization Benchmark [589], SCitance [17], and MissciPlus [243] are designed to complete scientific evidence collection. TheoremExplainBench [392] , XClaimCheck [365], SciNews [91], ClaimReview2024+ [75], FactKG [378], TrendFact [908], and SciVer [776] provide datasets and benchmarks for scientific verification analysis. MiniF2F [935], FIMO [482], MUSTARDSAUCE [325] are used to fine-tune and evaluate LLMs on scientific theorem proving tasks. Furthermore, datasets and benchmarks have expanded into broader domains, including biomedicine [814].
@@
To evaluate model performance in realistic research environments, MLAgentBench [322], Exp-Bench [385], MLRC-Bench [919], MLE-Bench [95], DS-Bench [352], ScienceBoard [730], ScienceArena [932], AutoReproduce [929], SciReplicate-Bench [844], DO Challenge [700], and MLR-Bench [103] assess AI agents’ abilities to perform typical research tasks, such as optimizing CIFAR-10 classifiers and tuning BabyLM. In a similar vein, Hu et al. [309] develop InfiAgent-DABench, which is based on real-world CSV datasets and evaluates models’ ability to interact with tools in end-to-end data analysis tasks. MLGym-Bench [561] is the first Gym environment for machine learning tasks, enabling research on reinforcement learning algorithms for training such agents. ResearchCodeBench [312] enables continuous understanding and advancement of LLM-driven innovation in research code generation. AutoBio [412] is designed to evaluate robotic automation in biology laboratory environments.
@@ 9.4 AI for Academic Writing
The field of AI in academic writing is supported by a comprehensive array of meticulously curated datasets that address various aspects of the academic writing process. Detailed tools are summarized in Table 9, which provides an overview of representative AI systems and their contributions to enhancing academic writing.
@@ 9.4.1 Semi-Automatic Academic Writing
Several datasets, including FigGen [646], SridBench [97], Figuring out Figures [90], SciCapenter [302], and TikZero [55], support figure and formula generation, from text-to-figure creation to automated TikZ code generation. For citation management, datasets like CITEWORTH [827], CiteBART [92], and ScholarCopilot [813] enhance context-aware automatic citation generation [107]. Additionally, FutureGen [35] extracts future work statements from thousands of papers, using LLMs to identify and validate forward-looking scientific content.
@@ 9.5 AI for Academic Peer Reviewing
Research on AI for Academic Peer Reviewing is grounded in diverse datasets, addressing tasks from AI text detection to review generation, quality assessment, and decision support [944, 197, 635, 139, 138, 715, 48]. To simulate realistic peer review interactions, datasets such as ReviewCritique [180], PeerRead [360], SPOT [703], NLPeer [183], ReviewMT [742], MOPRD [472], OpenReviewer [329], MASSW [909], COMPARE [693] , PeerArg [723], Re2 [897], ReviewEval [382], AAAR-1 [507], Papereval [323], ORB [736] and ORSUM [893] collect extensive paper and review data from leading conferences or journals, enabling the training and evaluation of LLMs in multi-turn, long-context, or role-based peer reviews [601]. Additionally, LLMart [463] offers a toolkit for evaluating LLM robustness through adversarial testing and prompt optimization, ensuring AI reliability in sensitive academic contexts. To assess LLM review quality more precisely, Shin et al. [682] and Couto et al. [149] analyze the quality of peer review content across multiple predefined aspects, highlighting discrepancies between LLM and human review focus.
@@ 10.4 Explainability and Transparency of AI4Research
As AI models increasingly drive scientific discovery, ensuring their trustworthiness, transparency and explainability is essential. Future work should strengthen model interpretability so that researchers can trace how conclusions and recommendations are generated, particularly in high-stakes scientific applications [214, 188, 802]. Efforts to improve explainability fall into two main categories: (1) White-box Analysis: This approach investigates the model’s internal structure by linking specific network “circuits” to conceptual representations. It has attracted considerable interest from both the security and transparency communities [57, 924, 630]. (2) Black-box Analysis: More recent work focuses on interpreting models without direct access to internal parameters. By examining reasoning trajectories and aggregate behavior, black-box methods provide insights into a model’s knowledge representation and enable more reliable control over its outputs [85, 282, 108, 110, 112].